OpenStudy (anonymous):

plsease help me solve this 3 (2x-4) =6+2 (x+5)

OpenStudy (anonymous):

Use distributive property to solve this problem. 6x-12=6 + 2x+ 10 6x-12= 16+ 2x 4x=28 x=7

OpenStudy (anonymous):

pls solve this for me 3x-4=4(x+3) pls show steps